OpenStudy (anonymous):

0.01x - 0.3y = 1 y = 0.1x - 2 What is the solution of the given system? (-20, -4) (0.4, 2) (0, 20) GOT B

OpenStudy (anonymous):

MEANT A

OpenStudy (anonymous):

?

OpenStudy (ikram002p):

\[\left[ 0.01x-0.3y=1 \right]*10\rightarrow 0.1x-3y=10 \] \[-2y=8,y=-4\rightarrow Apply y=-4 then you would get x=-20\] so you answer A is correct